• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 23:11:32 on October24,2009

    Entry number 296507

    keyword=Compton rates ok at 80 uA - DAQ ok

    I power cycled the ETROC and leave it with the old firmware until we take a coincidence run since this one does not have inhibit. Photon spectrum for photon looks ok

    FIGURE 1

    FIGURE 2